OpenStudy (anonymous):

Q6 Vin=15V, R1=0.5Ω, R2=1.5Ω, R3=5Ω, A=2 and B=1.5 Rth=?

OpenStudy (anonymous):

Vin = R1 Isc + R2 Isc IsC =Vin/ R1 + R2 Rth = Vth / Isc ?????

OpenStudy (anonymous):

vth = 3 vo = 3 I =2.4 Rth = 4

OpenStudy (nurali):

Question 6 Formula Vout = Vin*(R3-B*R1*R3)/(R3-(B*R1*R3)+R1+R2+A*R2) i1=V0/(R3-B*R1*R3) and Vth is same as Vout. Isc=vIN*(1-B*R1)/(A*R2+R1+R2) Rth=VO/iSC where VO is Vth

OpenStudy (nurali):

If correct please click best response

OpenStudy (anonymous):

can you please explain the steps a bit. I dont know how to treat Ava and Bvb

OpenStudy (anonymous):

what is the formula for vth when dependent sources are present.?